OpenStudy (anonymous):

Find the solution to x2 = 121 Show all your steps.

OpenStudy (anonymous):

x^2=11*11=11^2

OpenStudy (anonymous):

\[\sqrt{x ^{2}}=\sqrt{121}\]\[x=11\]